### v1

# Old Meets New: Connecting Two Infinite Families of Congruences Modulo Powers of 5 for Generalized Frobenius Partition Functions

Abstract.
In 2012 Paule and Radu proved a difficult family of congruences modulo powers of 5 for Andrews’ 2-colored generalized Frobenius partition function. The family is associated with the classical modular curve of level 20. We demonstrate the existence of a second congruence family for a related generalized Frobenius partition function associated with the same curve. Unlike traditional proofs, we construct an isomorphism between this new family and the original family of congruences via an automorphism on the associated ring of functions. We also give some important insights into the behavior of these congruence families with respect to the Atkin–Lehner involution which proved very important in Paule and Radu’s original proof.
Key words and phrases:
Partition congruences, infinite congruence family, modular functions, plane partitions, partition analysis, modular curve, Riemann surface
2010 Mathematics Subject Classification:
Primary 11P83, Secondary 30F35
1. Introduction
1.1. Background
In his 1984 AMS Memoir, Andrews [1] introduced and extensively studied two families of generalized Frobenius partition functions. The second such family of functions enumerates generalized $k$-colored Frobenius partitions of weight $n$, where the functions in question are denoted $c\phi_{k}(n)$. In his Memoir, Andrews proved a variety of properties satisfied by these functions $c\phi_{k}(n)$; in particular, Andrews proved that, for all $n\geq 0$, $c\phi_{2}(5n+3)\equiv 0\pmod{5}$.
In 1994 the first author proposed the following congruence family for the generalized 2-colored Frobenius partition function $c\phi_{2}(n)$:
Theorem 1.1 (Paule, Radu).
Let $n,\alpha\in\mathbb{Z}_{\geq 1}$ such that $12n\equiv 1\pmod{5^{\alpha}}$. Then $c\phi_{2}(n)\equiv 0\pmod{5^{\alpha}}$.
What makes this result so interesting is that, while it superficially resembles the classical congruence families of the unrestricted partition function $p(n)$ modulo powers of 5, 7, 11 that Ramanujan first discovered [10], it is a much more difficult result to prove. Indeed, a proof was not found until Paule and Radu’s work in 2012 [7], 18 years after its proposal, and some 93 years after Ramanujan founded the subject.
This family poses problems for us still. Like that of other congruence families, the proof given by Paule and Radu involves the $\ell$-adic properties of an associated sequence of meromorphic functions on a classical modular curve. However, the proof that Paule and Radu provide invokes a remarkable algebraic machinery that is not a priori obvious or natural in the framework of the corresponding modular curve.
The second author has invented some more general techniques, embodied in the localization method, in an attempt to give a unified framework to the subject of partition congruences. These techniques work in many different cases (e.g., [11], [12]), and indeed reduce to standard methods for classical congruence families. However, the congruences of Theorem 1.1 are strangely resistant to localization. As such, we cannot at present extend classical methods via a systematic approach to congruence families of a certain class which encompass Theorem 1.1 and other contemporary congruence families, some of which are yet standing conjectures. In a sense, we know that the family is correct, but we still do not fully understand why it is there, and why more systematic approaches fail to resolve it.
In this work we report on recent progress in studying the underlying causes of this family. To our surprise, we have discovered that Theorem 1.1 is but part of a natural and extremely closely related pair of congruence families, exhibited by a pair of Frobenius partition functions. It may be that one reason that the congruences for $c\phi_{2}$ have proved so difficult to understand is that in the decade since they were resolved (and the three decades since they were first proposed), they have been effectively half perceived.
This discovery arose after we became aware of a study of generalized Frobenius partition functions by Jiang, Rolen, and Woodbury [3]. They define a class of functions of the form $c\psi_{k,\beta}(n)$ for which
$c\psi_{2,1}(n)$ $=c\phi_{2}(n)$ (1.1)
for all $n\geq 0$. Given the precise definition of this class of functions, as well as the restrictions on $k$ and $\beta$ (see below), it is natural to consider the function $c\psi_{2,0}(n)$. Indeed, this function had been studied previously by Drake [5], who used the notation $c\psi_{2}(n)$. We discovered that $c\psi_{2}(n)$ exhibits the following congruence family:
Theorem 1.2.
Let $n,\alpha\in\mathbb{Z}_{\geq 1}$ such that $6n\equiv-1\pmod{5^{\alpha}}$. Then $c\psi_{2}(n)\equiv 0\pmod{5^{\alpha}}$.
Henceforth we will use the notation developed by Jiang, Rolen, and Woodbury [3], since it has proved so insightful; in particular, it allows us to express Theorems 1.1 and 1.2 together in the following form:
Theorem 1.3.
Fix $\beta\in\{0,1\}$ and let $n,\alpha\in\mathbb{Z}_{\geq 1}$ such that $3\cdot 2^{\beta+1}n\equiv(-1)^{\beta+1}\pmod{5^{\alpha}}$. Then
$c\psi_{2,\beta}(n)\equiv 0\pmod{5^{\alpha}}.$
The latter theorem already suggests a close relationship between the two families. However, we were surprised to discover just how close this connection is. In the first place, each family is associated with a sequence of functions defined on the classical modular curve $\mathrm{X}_{0}(20)$—a curve of genus 1 and cusp count 6. For each case, one constructs a pair of rank 2 $\mathbb{Z}[t]$-modules (with $t$ a natural Hauptmodul corresponding to $\mathrm{X}_{0}(5)$), and demonstrates membership of the respective functions within these modules.
For each family, the generators of the respective modules are different; however, we have found that the proof of Theorem 1.2 is—at least in terms of formal algebraic manipulations—identical to that of Theorem 1.1. Indeed, both congruence families are related to each other via a certain automorphism between subrings of the modular functions on $\mathrm{X}_{0}(20)$ which fixes the modular functions on $\mathrm{X}_{0}(5)$. This may have enormous implications on the future study of congruence families associated with modular curves of composite level.
We discovered a second strange property which emerges when we examine the module generators associated with each respective congruence family. Apart, each is a function defined on $\mathrm{X}_{0}(20)$. When combined, they produce a function defined on $\mathrm{X}_{0}(10)$, a curve of genus 0 and cusp count 4. This is especially interesting, in that congruence families associated with such curves, i.e., Riemann surfaces of simpler topologies, are much more accessible to our standardized proof methods.
Finally, these families bear a strange symmetry to one another upon the application of a certain Atkin–Lehner involution to their respective functions. This symmetry gives rise to an absolutely astonishing and beautiful result which explains a curious coefficient behavior that was discovered by Paule and Radu in their original work [7, Section 6], and used to assist in the construction of their proof of Theorem 1.1.
The remainder of our paper is organized as follows: in Section 2 we give the definitions for the associated Frobenius partitions as well as the generating functions relevant to our work. In Section 3 we give the underlying modular function spaces over $\mathrm{X}_{0}(20)$ that will be relevant to us. We prove the initial stages of both congruence families via the associated initial terms on the related function sequences. We then show how Theorem 1.2 can be derived from Theorem 1.1 via a certain automorphism on a subring of modular functions of $\mathrm{X}_{0}(20)$.
In Section 4 we show how these two sequences, when combined, produce functions which reduce to the simpler curve $\mathrm{X}_{0}(10)$. We then give conjectures as to the accessibility of proving the two congruence families relative to one another via more systematic techniques.
Finally, in Section 5 we show how a peculiar coefficient behavior discovered by Paule and Radu is the consequence of a very striking relationship with Ramanujan’s $\varphi$ function (i.e., Jacobi’s classic theta function $\vartheta_{3}(0|2\tau)$). In Section 6 we summarize and present the questions and conjectures which we consider especially important for further work.
In this work we presume an understanding of the theory of modular functions and its associated modular curves. See [11, Section 2]. A useful classical reference is [6], and a more modern reference is [4, Chapters 2-3].
[…]
2. Frobenius Partition Functions
As noted above, the general study of Frobenius partition functions was initiated by Andrews in 1984 [1]. This is now a broad area of study, and we will concentrate on a recent synthesis by Jiang, Rolen, and Woodbury [3] under which the binary nature of the Andrews–Sellers congruences can be recognized.
Definition 2.1.
(Jiang–Rolen–Woodbury) Let $n,k\in\mathbb{Z}_{\geq 1}$ and $\beta\in\mathbb{Z}+\frac{k}{2}$ nonnegative. A $(k,\beta)$-colored generalized Frobenius partition of $n$ is an array of the form
$\begin{pmatrix}a_{1}&a_{2}&...&a_{r}\\ b_{1}&b_{2}&...&b_{s}\end{pmatrix}$
satisfying the following:
  • Each $a_{i},b_{j}\in\mathbb{Z}_{\geq 0}$, and any integer occurs at most $k$ times,
  • Each row is decreasing with respect to lexicographic ordering,
  • $r+s\neq 0$
  • $r-s=\beta-\frac{k}{2}$
  • $n=r+\sum_{0\leq i\leq r}a_{i}+\sum_{0\leq j\leq s}b_{j}.$
The number of such arrays is denoted $c\psi_{k,\beta}(n)$, and the associated generating function over the variable $q$ is denoted $\mathrm{C}\Psi_{k,\beta}(q)$.
By and large, the generating functions for $\mathrm{C}\Psi_{k,\beta}(q)$ are not a priori obvious; indeed, a significant portion of [3] is devoted specifically to the problem of constructing such generating functions. In the cases $k=2$, $\beta=0,1$, the generating functions are simple to state [1, 5, 3]:
We have
$\mathrm{C}\Psi_{2,1}(q):=\sum_{n=0}^{\infty}c\phi_{2}(n)q^{n}$ $=\sum_{n=0}^{\infty}c\psi_{2,1}(n)q^{n}=\frac{(q^{2};q^{2})_{\infty}^{5}}{(q;q)_{\infty}^{4}(q^{4};q^{4})_{\infty}^{2}},$
$\mathrm{C}\Psi_{2,0}(q):=\sum_{n=0}^{\infty}c\psi_{2}(n)q^{n}$ $=\sum_{n=0}^{\infty}c\psi_{2,0}(n)q^{n}=\frac{(q^{4};q^{4})_{\infty}^{2}}{(q;q)_{\infty}^{2}(q^{2};q^{2})_{\infty}}.$
The close relationship between these two Frobenius functions is not clear from the generating functions themselves. On the other hand, Definition 2.1 reveals that the two are combinatorically close. The key difference is that for $c\psi_{2,1}(n)$ we must count arrays in which $r=s$; for $c\psi_{2,0}(n)$, we must have $s=r+1$.
3. Proof of Theorem 1.2
The generally accepted approach to proving a congruence family corresponding to the coefficients of a given modular form is to construct a sequence $\mathcal{L}:=(L_{\alpha})_{\alpha\geq 1}$ of functions on an associated manifold—which for our purposes will always be the classical modular curve $\mathrm{X}_{0}(N)$ for some positive integer $N$. Indeed, except for Section 4, we will always work with $\mathrm{X}_{0}(20)$. Each $L_{\alpha}$ enumerates a given case of the congruence family, and each function is related to its predecessor via a certain linear operator. The difficulty then is to express each $L_{\alpha}$ in terms of some convenient reference functions.
We will define a pair of function sequences,
$\mathcal{L}^{(\beta)}:=\left(L_{\alpha}^{(\beta)}\right)_{\alpha\geq 1}$ (3.1)
with $\beta\in\{0,1\}$, by
$L_{\alpha}^{(\beta)}:=\frac{1}{\mathrm{C}\Psi_{2,\beta}\left(q^{1+4a}\right)}\cdot\sum_{3\cdot 2^{\beta+1}n\equiv(-1)^{\beta+1}\bmod{5^{\alpha}}}c\psi_{2,\beta}(n)q^{\left\lfloor n/5^{\alpha}\right\rfloor+\beta},$ (3.2)
where $a\in\{0,1\}$, $\alpha\equiv a\pmod{2}$.
To relate each $L_{\alpha}^{(\beta)}$ to its successor, we construct a set of linear operators $U^{(\beta)}_{a}$. In particular, for $\beta=0,1$, we have
$U^{(\beta)}_{1}\left(\sum_{n\geq N}a(n)q^{n}\right)=\sum_{5n\geq N}a(5n)q^{n}$ (3.3)
is the standard $U_{5}$ operator (see [6, Chapter 8] for a list of the standard properties of $U_{5}$), and
$U^{(\beta)}_{0}\left(f\right)=U_{5}\left(\mathcal{A}^{(\beta)}\cdot f\right),$ (3.4)
where
$\mathcal{A}^{(\beta)}=q^{2(3\beta-2)}\frac{\mathrm{C}\Psi_{2,\beta}(q)}{\mathrm{C}\Psi_{2,\beta}(q^{25})}.$ (3.5)
It is a comparatively standard procedure in the theory to show that for all $\alpha\geq 1$ we have
$U^{(\beta)}_{a}\left(L_{\alpha}^{(\beta)}\right)=L_{\alpha+1}^{(\beta)}.$ (3.6)
See, for example, [2, p. 23], or [11, Lemma 2.13].
If we can show that $L_{\alpha}^{(\beta)}$ is divisible by $5^{\alpha}$, then we have proved Theorems 1.1 and 1.2. If we make the substitution $q=e^{2\pi i\tau}$ with $\tau\in\mathbb{H}$, then these function sequences are modular functions over the congruence subgroup $\Gamma_{0}(20)$; equivalently, they correspond to functions on the modular curve $\mathrm{X}_{0}(20)$.
This modular curve is a compact Riemann surface of cusp count 6 and genus 1. If we were to consider, say, the space of functions on $\mathrm{X}_{0}(20)$ which live at the cusp $[0]$ (that is, the meromorphic functions which are holomorphic along the entire curve except the cusp $[0]$), and denote this space as $\mathcal{M}^{0}\left(\mathrm{X}_{0}(20)\right)$, then the Weierstraß gap theorem (e.g., [8]) predicts that this space is a rank 2 $\mathbb{C}[x]$-module for some function $x$—that is, there exist two functions, which we will here denote $x$ and $y$, such that
$\mathcal{M}^{0}\left(\mathrm{X}_{0}(20)\right)=\mathbb{Z}[x]+y\mathbb{Z}[x].$ (3.7)
In this case, one may show with some ease that a sufficient choice of $x$, $y$ is
$x$ $=q\frac{(q^{2};q^{2})_{\infty}(q^{10};q^{10})_{\infty}^{3}}{(q;q)_{\infty}^{3}(q^{5};q^{5})_{\infty}},$ (3.8)
$y$ $=q^{2}\frac{(q^{2};q^{2})_{\infty}^{2}(q^{4};q^{4})_{\infty}(q^{5};q^{5})_{\infty}(q^{20};q^{20})_{\infty}^{3}}{(q;q)_{\infty}^{5}(q^{10};q^{10})_{\infty}^{2}}.$ (3.9)
Because the functions $L_{\alpha}^{(\beta)}$ are modular, they are equivalently meromorphic functions on $\mathrm{X}_{0}(20)$ with possible poles only at the cusps.
One of the features which substantially simplifies the proofs of other congruence families is that when the cusp count is 2, the associated functions $L_{\alpha}$ can be easily made to have positive order at one cusp. This forces them to have a pole at the other cusp, and consequently to be representable as a polynomial in the generators of the functions which live at that cusp. This is true, for example, in the cases of Ramanujan’s classic congruence families modulo powers of 5, 7, 11 (resp.) for the unrestricted partition function $p(n)$, in which the associated modular curves are $\mathrm{X}_{0}(\ell)$ with $\ell$ equal to 5, 7, 11 (resp.) [14, 2].
Notice that this is not possible in our current case. The cusp count is 6, and there is no guarantee that $L_{\alpha}^{(\beta)}$ live at the cusp $[0]$ (and indeed, they do not). One possible resolution is to multiply each $L_{\alpha}^{(\beta)}$ by an eta quotient which lives at $[0]$ and which has positive order high enough to eliminate all of the poles of $L_{\alpha}^{(\beta)}$ except for one at $[0]$. For us, such a function arises as powers of $1+5x$. This places $L_{\alpha}^{(\beta)}$ in a localized ring of modular functions. For example,
$L_{1}^{(1)}=\frac{5}{(1+5x)^{2}}$ $\bigg{(}4x+137x^{2}+1704x^{3}+10080x^{4}+28800x^{5}+32000x^{6}$
$-y(20+400x+3040x^{2}+10240x^{3}+12800x^{4})\bigg{)}.$ (3.10)
We instantly recognize divisibility of the function by 5 in this identity. Similar identities can easily be computed for larger $\alpha$, and it can be shown with relative ease that
$L_{\alpha}^{(1)}\in\mathbb{Z}[x]_{\mathcal{S}}+y\mathbb{Z}[x]_{\mathcal{S}},$ (3.11)
for
$\mathcal{S}:=\left\{(1+5x)^{n}:n\geq 0\right\}.$ (3.12)
The possibility of actually proving the whole family by manipulation of these rational polynomials is naturally very alluring. Such a class of proofs does exist, and has been exhibited for other difficult congruence families in previous papers.
This technique fails spectacularly in the cases of the families in Theorems 1.1 and 1.2. It is certainly true that manipulation of these rational polynomial identities allows one to confirm as many cases of either theorem as we like. However, for extremely deep and technical reasons, it is not currently possible to properly complete the induction along the lines of localization.
Paule and Radu concentrated on a different approach in [7]. They tried to express $L_{\alpha}^{(\beta)}$ almost entirely in terms of functions on a modular curve of lower level. Let us consider
$t=$ $\ \frac{\eta(5\tau)^{6}}{\eta(\tau)^{6}}=q\left(\frac{(q^{5};q^{5})_{\infty}}{(q;q)_{\infty}}\right)^{6}.$ (3.13)
This is the Hauptmodul for $\mathrm{X}_{0}(5)$, a curve of genus 0 and cusp count 2. Indeed, $t$ is the critical function for proving Ramanujan’s congruence family for $p(n)$ modulo powers of 5 (see [14] or [6, Chapters 7-8]). Being modular over $\Gamma_{0}(5)$, it is clearly modular over $\Gamma_{0}(20)$. It does not live at the cusp $[0]$, but we can perform the same trick we did with $L_{1}^{(1)}$—we can represent $t$ as a rational polynomial. Indeed, we have
$t=\frac{x(1+4x)^{2}}{1+5x}$ (3.14)
(this is proved by a straightforward procedure in our Mathematica supplement, which may be found online in the url at the end of Section 1). Notice that if we multiply through by $1+5x$, substitute $z=1+5x$, and clear fractions, we then have
$(1+5x)t$ $=x(1+4x)^{2}$ (3.15)
$zt$ $=\frac{-1}{125}(1+7z+8z^{2}-16z^{3}),$ (3.16)
$-125zt$ $=1+7z+8z^{2}-16z^{3}.$ (3.17)
If we then isolate 1, divide by $z$, and shift back to $x$, we of course have
11 11 $=-125tz-7z-8z^{2}+16z^{3},$ (3.18)
$\frac{1}{z}$ $=-125t-7-8z+16z^{2},$ (3.19)
$\frac{1}{1+5x}$ $=1-125t+120x+400x^{2}.$ (3.20)
Thus through an elementary set of manipulations, we can now represent elements of our localization ring as polynomials in $t$, $x$, $y$. Indeed, one sees that $x^{3}$ can be written as a polynomial in $t$, $x$, $x^{2}$. As such, we can express each $L_{\alpha}^{(\beta)}$ as a polynomial in $t,x,y$, but with the powers of $x$ restricted to 2 or lower (and the powers of $y$ restricted to 1 or lower). Thus, our representations which will be “mostly” in terms of the function $t$.
Nevertheless, from these manipulations, we would guess that the associated polynomials would occur as members of a rank 6 $\mathbb{Z}[t]$-module (corresponding to the basis elements 1, $x$, $x^{2}$, $y$, $yx$, $yx^{2}$). The second (and most important) breakthrough of Paule and Radu was to recognize that each $L_{\alpha}^{(1)}$ is expressible in terms of a $\mathbb{Z}[t]$-module of rank 3. Indeed, each $L_{\alpha}^{(1)}$ can be placed in one of a pair of rank 2 $\mathbb{Z}[t]$-modules depending on the parity of $\alpha$; that is,
$L_{\alpha}^{(1)}\in\mathbb{Z}[t]+p^{(1)}_{a}\mathbb{Z}[t]$ (3.21)
for a pair of functions $p^{(1)}_{0}$, $p^{(1)}_{1}$. It is the set of coefficients in this representation which may be closely studied to prove divisibility by ever increasing powers of 5.
For instance,
$L_{1}^{(1)}=-5t+25p^{(1)}_{1}.$ (3.22)
The question of how to derive this reduced rank representation is a serious one; for the time being we postpone discussion of this to Section 5, and instead simply give the functions in terms of $1,x,x^{2},y,xy,x^{2}y$, and powers of $t$:
$p^{(1)}_{0}=$ $\ 85t+625t^{2}-24x-64x^{2}-4y-1000ty-12500t^{2}y+992xy+12000txy+3520x^{2}y+40000tx^{2}y,$ (3.23)
$p^{(1)}_{1}=$ $\ t+25t^{2}-8tx-4y-8ty-500t^{2}y-32xy+480txy-64x^{2}y+1600tx^{2}y.$ (3.24)
Upon discovering the congruence family of Theorem 1.2, we guessed that the associated functions $L_{\alpha}^{(0)}$ could be represented in a similar manner. We eventually discovered that
$L_{\alpha}^{(0)}\in\mathbb{Z}[t]+p^{(0)}_{a}\mathbb{Z}[t],$ (3.25)
with
$p^{(0)}_{0}=$ $\ 1+5t+64x+600tx+208x^{2}+2000tx^{2}+4y+1000ty+12500t^{2}y-992xy-12000txy$
$-3520x^{2}y-40000tx^{2}y,$ (3.26)
$p^{(0)}_{1}=$ $\ 1+9t+8x+64tx+16x^{2}+80tx^{2}+4y+8ty+500t^{2}y+32xy-480txy+64x^{2}y-1600tx^{2}y.$ (3.27)
This seemed reasonable enough; however, comparison of the actual representations of $L_{\alpha}^{(0)}$, $L_{\alpha}^{(1)}$ led us to a much more stunning result. Let us give $L_{1}^{(0)}$:
$L_{1}^{(0)}=-5t+25p^{(0)}_{1}.$ (3.28)
Comparing (3.22) and (3.28), one finds that the two identifies formally match. Indeed, we have verified that the proofs, as well as the representations themselves, formally match exactly—they differ only in the exact form of the module generators $p^{(\beta)}_{\alpha}$.
Theorem 3.1.
Define
$\mathcal{R}:=\mathbb{C}[t]+p^{(0)}_{0}\mathbb{C}[t]\ +$ $p^{(0)}_{1}\mathbb{C}[t]+p^{(1)}_{0}\mathbb{C}[t]+p^{(1)}_{1}\mathbb{C}[t],$ (3.29)
and the ring automorphisms
$\sigma^{(\beta)}$ $:\mathcal{R}\longrightarrow\mathcal{R}$ (3.30)
:$:\left\{\begin{matrix}&t\longmapsto t,\\ &p^{(\beta)}_{a}\longmapsto p^{(1-\beta)}_{a}.\end{matrix}\right.$ (3.33)
Then
$\sigma^{(\beta)}\left(L^{(\beta)}_{\alpha}\right)=L^{(1-\beta)}_{\alpha}.$ (3.34)
This theorem is proved by examining the action of $U^{(\beta)}_{a}$ on $t^{m}$ and $t^{m}p^{(\beta)}_{a}$. That is, one simply must show that $U^{(0)}_{a}\left(t^{m}\right)$ (resp. $U^{(0)}_{a}\left(t^{m}p^{(0)}_{a}\right)$) formally matches $U^{(1)}_{a}\left(t^{m}\right)$ (resp. $U^{(1)}_{a}\left(t^{m}p^{(1)}_{a}\right)$) for $a=0,1$; the only difference is switching out one generator ($p^{(0)}_{a}$) for another ($p^{(1)}_{a}$).
In particular, $t$ is subject to a modular equation of degree 5 which is not dependent on our module generators [7, Definition 3.1, Lemma 3.2]. With this equation we can build a recurrence of $U^{(\beta)}_{a}\left(t^{m}\right)$, $U^{(\beta)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for any $m$, provided that we have precise expressions for 5 consecutive values of $m$ [7, Lemma 3.3].
In our Mathematica supplement (see the end of Section 1) we derive the explicit relations of $U^{(0)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for $a=0,1$, and five consecutive values of $m$. We then show that they match the relations of $U^{(1)}_{a}\left(t^{m}p^{(0)}_{a}\right)$ for the same values of $a$, $m$, which are derived by Paule and Radu in [7, Section 7].
In our case, we use the same values of $a$, $m$ that Paule and Radu used to expedite the comparison. We can prove these relations with the modular cusp analysis. This is also done in our Mathematica supplement (see the end of Section 1).
Of course, Theorem 1.1 was proved by showing that $L^{(1)}_{\alpha}$ is divisible by $5^{\alpha}$. Therefore, $\sigma^{(1)}\left(L^{(1)}_{\alpha}\right)=L^{(0)}_{\alpha}$ will also be so divisible. With that we immediately have Theorems 1.2 and 1.3.
4. Level Reduction Theorem
Notice that $\sigma^{(\beta)}$ is a mapping on the ring of modular functions over $\Gamma_{0}(20)$ which fixes the functions over $\Gamma_{0}(5)$ (since $t$ is a Hauptmodul for functions over the latter). Thus, the two congruence families are closely related to functions on modular function fields. This very likely poses significant implications for the question of how congruence families relate through mappings of modular functions on various congruence subgroups. We also note another strange result, perhaps in a similar vein. On comparing (3.23)-(3.24) with (3.26)-(3.27), one discovers the following:
$p^{(0)}_{a}+p^{(1)}_{a}\in\mathbb{Z}[x]_{\mathcal{S}}.$ (4.1)
That is, the contribution in $y$ is annihilated. Moreover, notice that by our choice, $x$ is not only modular over $\Gamma_{0}(20)$, but indeed over $\Gamma_{0}(10)$. So not only does the addition of our module generators annihilate the $y$ terms, it actually reduces the functions entirely to a curve of lower level.
If we combine (4.1) with Theorem 3.1, and remember from (3.14) that $t$ can be expressed as a member of $\mathbb{Z}[x]_{\mathcal{S}}$, we immediately have the following:
Theorem 4.1.
Let
$L_{\alpha}:=L_{\alpha}^{(0)}+L_{\alpha}^{(1)},$ (4.2)
Then
$L_{\alpha}\in\mathcal{M}\left(\mathrm{X}_{0}(10)\right).$
and
$L_{\alpha}\equiv 0\pmod{5^{\alpha}}.$
While the methods of localization described above have failed to resolve Theorems 1.1 and 1.2, it has been used successfully on congruence families associated with $\mathrm{X}_{0}(10)$. As such, Theorem 4.1 implies that localization may be able to prove the truth of one congruence family relative to the other. That is, localization is not sufficient to prove Theorem 1.1; however, it may be able to prove Theorem 1.1 provided that Theorem 1.2 is true, and vice versa.
We are as of yet unable to immediately prove this application of localization, though we believe that it is accessible.
Conjecture 4.2.
Theorem 4.1 is accessible via the localization method.
5. The Atkin–Lehner Involution
We close our report with a discussion of the striking dual relationship our congruence pair has with respect to Ramanujan’s classic $\varphi(q)$ function via a certain Atkin–Lehner involution. This partly answers (or at least extends) some of the questions which Paule and Radu posed in their proof of Theorem 1.1. We consider the Atkin–Lehner involution arising from the action of the matrix
$W:=\begin{pmatrix}4&-1\\ 100&-24\end{pmatrix}$
on our variable $\tau\in\mathbb{H}$.
Theorem 5.1.
Let $\text{Let }f=y^{j}\frac{x^{m}}{(1+5x)^{n}}\in\mathbb{Z}[x]_{\mathcal{S}}+y\mathbb{Z}[x]_{\mathcal{S}}\text{, for $j\in\{0,1\},$ and $m,n\in\mathbb{Z}_{\geq 0}$. Then }$
$f(W\tau)=(-1)^{m+n}4^{n-2j-m}\frac{\left(1+4x\right)^{j}\left(1+10x+20y\right)^{j}\left(1+10x+20x^{2}+4y\right)^{m}\left(1+5x+5y\right)^{n}}{(1+5x)^{3j+2m+n}}.$
Proof.
The functions $x$ and $y$ are modular over $\Gamma_{0}(20)$. Each is itself an eta quotient, whose forms we repeat here:
$x$ $=q\frac{(q^{2};q^{2})_{\infty}(q^{10};q^{10})_{\infty}^{3}}{(q;q)_{\infty}^{3}(q^{5};q^{5})_{\infty}}=\frac{\eta(2\tau)\eta(10\tau)^{3}}{\eta(\tau)^{3}\eta(5\tau)},$
$y$ $=q^{2}\frac{(q^{2};q^{2})_{\infty}^{2}(q^{4};q^{4})_{\infty}(q^{5};q^{5})_{\infty}(q^{20};q^{20})_{\infty}^{3}}{(q;q)_{\infty}^{5}(q^{10};q^{10})_{\infty}^{2}}=\frac{\eta(2\tau)^{2}\eta(4\tau)\eta(5\tau)\eta(20\tau)^{3}}{\eta(\tau)^{5}\eta(10\tau)^{2}}.$
Also note that $1+5x$ is an eta quotient, since we have
$1+5x=\frac{(q^{2};q^{2})_{\infty}^{5}(q^{5};q^{5})_{\infty}}{(q;q)_{\infty}^{5}(q^{10};q^{10})_{\infty}}=\frac{\eta(2\tau)^{5}\eta(5\tau)}{\eta(\tau)^{5}\eta(10\tau)}.$ (5.1)
Therefore, consider the action of $W$ on $\eta\left(\delta\tau\right)$ for $\delta|20$. We have
$\delta W(\tau)=\frac{4\delta\tau-\delta}{100\tau-24}=\begin{pmatrix}4\delta&-\delta\\ 100&-24\end{pmatrix}(\tau).$ (5.4)
We want to express this in terms of the action of $\mathrm{SL}\left(2,\mathbb{Z}\right)$. For example, we have
$W(\tau)$ $=\frac{4\tau-1}{100\tau-24}=\begin{pmatrix}1&0\\ 25&1\end{pmatrix}(4\tau-1),$ (5.7)
$2W(\tau)$ $=\frac{8\tau-2}{100\tau-24}=\begin{pmatrix}2&1\\ 25&13\end{pmatrix}(2\tau-1),$ (5.10)
$4W(\tau)$ $=\frac{16\tau-4}{100\tau-24}=\begin{pmatrix}4&-1\\ 25&-6\end{pmatrix}(\tau),$ (5.13)
$5W(\tau)$ $=\frac{20\tau-5}{100\tau-24}=\begin{pmatrix}1&0\\ 5&1\end{pmatrix}(20\tau-5),$ (5.16)
$10W(\tau)$ $=\frac{40\tau-10}{100\tau-24}=\begin{pmatrix}2&1\\ 5&3\end{pmatrix}(10\tau-3),$ (5.19)
$20W(\tau)$ $=\frac{80\tau-20}{100\tau-24}=\begin{pmatrix}4&-1\\ 5&-1\end{pmatrix}(5\tau-1).$ (5.22)
If we consider, say, the involution applied to $x$, we have
$x\left(W\tau\right)$ $=\frac{\eta(2W\tau)\eta(10W\tau)^{3}}{\eta(W\tau)^{3}\eta(5W\tau)}$ (5.23)
$=\frac{\eta\left(\begin{pmatrix}2&1\\ 25&13\end{pmatrix}(2\tau-1)\right)\eta\left(\begin{pmatrix}2&1\\ 5&3\end{pmatrix}(10\tau-3)\right)^{3}}{\eta\left(\begin{pmatrix}1&0\\ 25&1\end{pmatrix}(4\tau-1)\right)^{3}\eta\left(\begin{pmatrix}1&0\\ 5&1\end{pmatrix}(20\tau-5)\right)}$ (5.32)
$=\frac{\left(-i\left(50\tau-12\right)\right)^{1/2}\left(-i\left(50\tau-12\right)\right)^{3/2}}{\left(-i\left(100\tau-24\right)\right)^{3/2}\left(-i\left(100\tau-24\right)\right)^{1/2}}\cdot\frac{\epsilon(2,1,25,13)\epsilon(2,1,5,3)^{3}}{\epsilon(1,0,25,1)^{3}\epsilon(1,0,5,1)}\cdot\frac{\eta(2\tau-1)\eta(10\tau-3)^{3}}{\eta(4\tau-1)^{3}\eta(20\tau-5)}$ (5.33)
$=\frac{1}{4}\frac{\epsilon(2,1,25,13)\epsilon(2,1,5,3)^{3}}{\epsilon(1,0,25,1)^{3}\epsilon(1,0,5,1)}\cdot\exp\left(\frac{2\pi i}{12}\left(2\tau-1+3(10\tau-3)-3(4\tau-1)-20\tau+5\right)\right)$ (5.34)
×$\times\frac{(q^{2};q^{2})_{\infty}(q^{10};q^{10})_{\infty}^{3}}{(q^{4};q^{4})_{\infty}^{3}(q^{20};q^{20})_{\infty}}$ (5.35)
$=-\frac{1}{4}\frac{(q^{2};q^{2})_{\infty}(q^{10};q^{10})_{\infty}^{3}}{(q^{4};q^{4})_{\infty}^{3}(q^{20};q^{20})_{\infty}}.$ (5.36)
Notice that $x(W\tau)$ is modular over $\Gamma_{0}(20)$. We next note that
$(1+5x)^{2}x(W\tau)\in\mathcal{M}^{0}\left(\mathrm{X}_{0}(20)\right)=\mathbb{Z}[x]+y\mathbb{Z}[x].$ (5.37)
We can then compute that
$(1+5x)^{2}x(W\tau)=-(1+10x+20x^{2}+4y).$ (5.38)
In a similar manner, we can show that
$(1+5x)^{3}y(W\tau)=(1+4x)(1+10x+20y),$ (5.39)
$(1+5x)\cdot\frac{1}{1+5x(W\tau)}=(1+5x+5y).$ (5.40)
We thus have a means of studying the direct effect of $W$ on our function spaces. Now, Paule and Radu discovered that
$\frac{(q^{10};q^{10})^{5}_{\infty}}{(q^{5};q^{5})^{2}_{\infty}}L_{2\alpha-1}^{(1)}(W\tau)$ $=\sum_{n=0}^{\infty}c_{2\alpha-1}^{(1)}(4n)q^{4n}+q\sum_{n=0}^{\infty}c_{2\alpha-1}^{(1)}(4n+1)q^{4n},$ (5.41)
$\frac{(q^{2};q^{2})^{5}_{\infty}}{(q;q)^{2}_{\infty}}L_{2\alpha}^{(1)}(W\tau)$ $=\sum_{n=0}^{\infty}c_{2\alpha}^{(1)}(4n)q^{4n}+q\sum_{n=0}^{\infty}c_{2\alpha}^{(1)}(4n+1)q^{4n}.$ (5.42)
This played a vital role in recognition of the module generators $p^{(1)}_{0},p^{(1)}_{1}$.
Upon becoming aware of the congruence family of Theorem 1.2, we attempted to find similar properties for the functions $L_{\alpha}^{(0)}$. After some experimentation, which included derivation of Theorem 5.1 above, we discovered that
$\frac{(q^{5};q^{5})^{2}_{\infty}}{(q^{10};q^{10})_{\infty}}L_{2\alpha-1}^{(0)}(W\tau)$ $=\sum_{n=0}^{\infty}c_{2\alpha-1}^{(0)}(4n)q^{4n}+q\sum_{n=0}^{\infty}c_{2\alpha-1}^{(0)}(4n+1)q^{4n},$ (5.43)
$\frac{(q;q)^{2}_{\infty}}{(q^{2};q^{2})_{\infty}}L_{2\alpha}^{(0)}(W\tau)$ $=\sum_{n=0}^{\infty}c_{2\alpha}^{(0)}(4n)q^{4n}+q\sum_{n=0}^{\infty}c_{2\alpha}^{(0)}(4n+1)q^{4n}.$ (5.44)
The first author of this paper recognized the pattern between these relationships. Notice that multiplication or division by any powers of $(q^{4};q^{4})_{\infty}$ will not affect the behavior of the powers of $q$ that Paule and Radu noticed. With this in mind, we recall Ramanujan’s phi function, effectively the theta function
$\varphi(q):=\frac{(q^{2};q^{2})^{5}_{\infty}}{(q;q)^{2}_{\infty}(q^{4};q^{4})_{\infty}^{2}}=\sum_{n=-\infty}^{\infty}q^{n^{2}}.$ (5.45)
Notice that
$\varphi(-q):=\frac{(q;q)^{2}_{\infty}}{(q^{2};q^{2})_{\infty}}.$ (5.46)
With these, we have the following result, beautiful and strange:
Theorem 5.2.
$\varphi\left((-1)^{\beta+1}q^{1+4a}\right)L_{\alpha}^{(\beta)}(W\tau)\in\mathbb{Z}[[q^{4}]]+q\mathbb{Z}[[q^{4}]].$ (5.47)
Proof.
This can be demonstrated by examining the $q$ subsequences in the progressions $24n+$ 2, $34n+$ 3 and showing that they vanish. To do this, we first note (as Paule and Radu did) that progressions other than $4n$ all vanish for $t(W\tau)$. We need only concern ourselves, then, with the effect of $W$ on $p^{(\beta)}_{a}$.
Beginning with $a=1$, we apply $W$ to $p^{(\beta)}_{1}$ by Theorem 5.1 above. This gives us a combination of eta quotients. We multiply each eta quotient by $\varphi\left((-1)^{\beta}q^{5}\right)$. We no longer have modular functions over $\Gamma_{0}(20)$, but that is immaterial to us. We next examine the progressions $34n+2,4n+$ 3 that each eta quotient will contribute (providing of course for the coefficients of each eta quotient and its leading power of $q$). Finally, we show that both progressions vanish. This is done in our Mathematica supplement (see the end of Section 1), via a careful manipulation of the Ramanujan–Kolberg algorithm which was originally designed by Radu [9] and implemented in [13].
In the case of even-indexed $\alpha$, we need only note that
$\varphi\left((-1)^{\beta+1}q\right)L_{2\alpha}^{(\beta)}(W\tau)$ $=U^{(\beta)}_{1}\left(\varphi\left((-1)^{\beta+1}q^{5}\right)L_{2\alpha-1}^{(\beta)}(W\tau)\right)$ (5.48)
$=U_{5}\left(\varphi\left((-1)^{\beta}q^{5}\right)L_{2\alpha-1}^{(\beta)}(W\tau)\right).$ (5.49)
Of course, $U_{5}$ extracts the coefficients of the form $5n$. But $5n\equiv 2,3\pmod{4}$ if and only if $n\equiv 2,3\pmod{4}$. Thus, these progressions must also vanish. ∎
It is certainly possible that a simpler proof may yet be found. Indeed, in a similar vein we produce the following conjecture:
Conjecture 5.3.
$\sum_{\beta=0,1}\varphi\left((-1)^{\beta+1}q^{1+4a}\right)L_{\alpha}^{(\beta)}(W\tau)\in\mathbb{Z}[[q^{4}]].$ (5.50)
This, we are certain, can be established by a similarly careful manipulation of the progressions for $14n+$ 1 in each of the series. We have only stopped for want of our time to other immediate demands.
6. Further Work
To the interested reader we emphasize five considerations which we deem important for future work.
  1. (1)
    Is our system of congruences for $c\psi_{2,\beta}(n)$ complete for $\beta=0,1$? That is to say, is this pair part of a broader triple or quadruple of families all closely related via their proofs or their behavior through some associated Atkin–Lehner involutions? Are there perhaps other mappings similar to $\sigma$ which would relate these families to others? Indeed, we believe that this specific pair of congruence families is not so closely related to any others in the same way that they are to each other—but we cannot be certain.
  2. (2)
    What are the implications of identifying different congruence families with one another via automorphisms which fix certain sub-module function spaces?
  3. (3)
    Do similar systems of congruence families exist for more general $c\psi_{k,\beta}(n)$, especially by holding a given $k$ fixed and varying $\beta$?
  4. (4)
    Is Conjecture 4.2 accessible? What implications does this pose for other congruence families, especially those associated with modular curves of large cusp count?
  5. (5)
    What are we to think of the beautiful yet utterly bizarre occurence of Ramanujan’s phi function in Theorem 5.2 which connects Theorems 1.1 and 1.2 via the chosen Atkin–Lehner involution? Can this be generalized to other congruence families, again with emphasis on those associated with curves of large cusp count? What of Conjecture 5.3?
7. Acknowledgments
The second author was funded in whole by the Austrian Science Fund (FWF) Principal Investigator Project 10.55776/PAT6428623, “Towards a Unified Theory of Partition Congruences.”